b1az3:

Solve the following system. y = -2x^2 + 3x - 1 and x + y = -1. The solutions are and

AZ:

How would you like to solve the system of equations? We can use the substitution method and solve a quadratics equation or we could graph the two equations and find their points of intersection. Do you know how to substitute the first equation into the second one? Or how to graph the equations to find where they intersect?

AZ:

For the substitution method, let me add some colors \( \color{red}{y} = \color{red}{-2x^2 + 3x - 1}\) and \( x + \color{red}{y} = -1\) since y = that long equation you can replace the y in the second equation with that long thing like \( x +( \color{red}{-2x^2 + 3x - 1}) = -1\) so now we have \( x -2x^2 + 3x - 1= -1\) can you simplify it and solve for x?

AZ:

If you want to graph it, you can use: https://www.desmos.com/calculator